# An Evaluation of Explanation Methods for Black-Box Detectors of Machine-Generated Text

### v1

@@ Appendix B Appendix: User Study
Note that these selection strategies will yield different sets for different pairings of detectors and explanation methods. The experiment are conducted with three sets of pairs, one per detector (Figure 4). Of the 18 pairs in a set, 6 will be determined by each explanation method. Document- and explanation similarity for the three sets are reported in Table 3. The datasets are balanced in respect to the detectors’ predictions, to eliminate counting examples in phases 1 or 3 to infer the number of machine- and human-documents in phases 2 or 4 as a strategy Hase and Bansal (2020).
Assignment
@@
Table 3 reports the document- and explanation similarity of pairs obtained with the proposed selection strategy. Pairs of explanations are significantly more similar than random selections, while not featuring overly similar documents.
((a)) Document Similarity
@@
((b)) Explanation Similarity
@@
Cosine Sim
FI-Features
Solaiman 0.31 0.17 0.15
@@
# Matching
Anchors
Solaiman 0.33 0.09 0.24
@@
Table 3: Similarity between pairs $(a_{j},b_{j})$ in the datasets for the user study against the mean similarity of 10 random selections (p < 0.05 bold). Cosine similarity is given as the mean of SHAP and LIME
Appendix C Appendix: Annotator Guidelines and Examples
For each document, please also rate to what extent you agree with these statements:
  • Q1
    From the explanation, I understand why the detector decided the way it did for this document.
    Select agree or strongly agree if you think the visualization presents sufficient evidence for why the detector decided the way it did in this specific case.
    Select disagree or strongly disagree if you can’t figure out why the detector decided the way it did.
  • Q2
    From the explanation, I now better understand how the detector works.
    Select agree or strongly agree if you think this explanation increased your understanding of how the detector reasons.
    Select disagree or strongly disagree if you don’t.
  • Q3
    The information from this explanation will help me predict the detector’s behaviour.
    Select agree or strongly agree if you think you could apply this information to the documents you labelled in the previous phase. You will do so in the next phase.
    Select disagree or strongly disagree if you don’t.
@@ Appendix C Appendix: Annotator Guidelines and Examples
Figure 6: Instructions on SHAP and LIME shown in Phase 3
@@
((a)) Set Guo, selected by SHAP; $y_{i}=f(d_{i})=\text{machine}$
((b)) Set DetectGPT, selected by LIME; $y_{i}=f(d_{i})=\text{human}$
((a)) Set Solaiman, selected by Anchor; $y_{i}=f(d_{i})=\text{machine}$
Figure 8: Pairs of documents shown to users. Users only see explanations for the first document of each pair and annotate the second one.
Appendix D Appendix: Complementary Results
Acc F1 AUC TN FP FN TP ms/evaluation
DetectGPT GPT-2 @100 samples 0.502 0.000 0.500 153 0 152 0 15808
DetectGPT pythia-70m @100 samples 0.705 0.579 0.704 153 0 90 62 6391
DetectGPT pythia-70m (this paper) 0.744 0.664 0.743 150 3 75 77 898
Solaiman 0.921 0.922 0.921 139 14 10 142 19
Guo 0.990 0.990 0.990 153 0 3 149 18
Table 4: Performance of the detectors. Note that one could obtain better results for DetectGPT when using GPT-2 by adjusting the classification threshold
DetectGPT00100100100100200200200200300300300300count f(x) = machine0%single5%10%25%50%80%masked00100100100100200200200200300300300300count f(x) = humanGuo00100100100100200200200200300300300300count f(x) = machine00100100100100200200200200300300300300count f(x) = humanSolaiman00100100100100200200200200300300300300count f(x) = machine00100100100100200200200200300300300300count f(x) = humanpadspaceremove
@@ Appendix D Appendix: Complementary Results
((a)) Guo
Explainer Score p
LIME 0.605 0.168
@@
((b)) Solaiman
Explainer Score p
LIME 0.402 0.004
@@
((c)) DetectGPT
Explainer Score p
Random 0.577 0.000
@@
((a)) Consistency
@@
((b)) Continuity
Explainer Detector $\alpha$
SHAP Guo 0.896
Solaiman 0.812
DetectGPT 0.081
LIME Guo 0.478
Solaiman 0.439
DetectGPT 0.265
Anchor Guo 0.367
Solaiman 0.135
DetectGPT 0.129
Random Solaiman -0.137
Guo -0.137
DetectGPT -0.144
Figure 10: Consistency and Continuity: Per detector results
@@
((b)) Fraction of tokens edited until $f(d_{i}^{\Omega})\neq f(d_{i})$
User Acc without User Acc with Change p
LIME Solaiman 0.681 0.569 -16.33% 0.057
LIME DetectGPT 0.792 0.681 -14.04% 0.134
LIME Guo 0.750 0.681 -9.26% 0.359
Anchor DetectGPT 0.806 0.750 -6.90% 0.289
SHAP DetectGPT 0.819 0.806 -1.69% 1.000
SHAP Solaiman 0.667 0.681 2.08% 1.000
Anchor Solaiman 0.569 0.583 2.44% 1.000
Anchor Guo 0.708 0.764 7.84% 0.344
SHAP Guo 0.778 0.847 8.93% 0.227
DetectGPT Anchor Q1 2.15 Q2 2.04 Q3 2.11 LIME Q1 3.62 Q2 2.97 Q3 3.08 SHAP Q1 2.54 Q2 2.56 Q3 2.57 Solaiman Anchor Q1 3.17 Q2 3.18 Q3 3.15 LIME Q1 4.00 Q2 3.81 Q3 3.65 SHAP Q1 3.33 Q2 3.39 Q3 3.29 Guo Anchor Q1 2.39 Q2 2.22 Q3 2.26 LIME Q1 3.17 Q2 3.32 Q3 3.19 SHAP Q1 3.29 Q2 2.64 Q3 2.65
Table 6: Forward Simulation and Perceived Usefulness: Per group results. (3 = neutral, 5 = strongly agree)

### v2

@@ Appendix A User study
Note that these selection strategies will yield different sets for different pairings of detectors and explanation methods. The experiments are conducted with three sets of pairs, one per detector (Figure 6). Of the 18 pairs in a set, 6 will be determined by each explanation method. Document- and explanation similarity for the three sets are reported in tables 3 and 4. The datasets are balanced in respect to the detectors’ predictions, to eliminate counting examples in phases 1 or 3 to infer the number of machine- and human-documents in phases 2 or 4 as a strategy Hase and Bansal (2020).
Assignment.
@@
Tables 3 and 4 report the document- and explanation similarity of pairs obtained with the proposed selection strategy. Pairs of explanations are significantly more similar than random selections, while not featuring overly similar documents.
@@
Table 3: Document Similarity between pairs $(a_{j},b_{j})$ in the datasets for the user study against the mean similarity of 10 random selections ($p<0.05$ bold). Cosine similarity is given as the mean of SHAP and LIME
@@
Cosine Sim FI-Features Solaiman 0.31 0.17 0.15
@@
# Matching Anchors Solaiman 0.33 0.09 0.24
@@
Table 4: Explanation Similarity between pairs $(a_{j},b_{j})$ in the datasets for the user study against the mean similarity of 10 random selections ($p<0.05$ bold). Cosine similarity is given as the mean of SHAP and LIME
For each document, please also rate to what extent you agree with these statements: Q1 From the explanation, I understand why the detector decided the way it did for this document. Select agree or strongly agree if you think the visualization presents sufficient evidence for why the detector decided the way it did in this specific case. Select disagree or strongly disagree if you can’t figure out why the detector decided the way it did. Q2 From the explanation, I now better understand how the detector works. Select agree or strongly agree if you think this explanation increased your understanding of how the detector reasons. Select disagree or strongly disagree if you don’t. Q3 The information from this explanation will help me predict the detector’s behaviour. Select agree or strongly agree if you think you could apply this information to the documents you labelled in the previous phase. You will do so in the next phase. Select disagree or strongly disagree if you don’t.
@@
Figure 8: Instructions on LIME shown in Phase 3
Figure 9: Instructions on SHAP shown in Phase 3
@@
Figure 11: Set Guo, selected by SHAP; $y_{i}=f(d_{i})=\text{machine}$. Users only see explanations for the first document of each pair and annotate the second one
Figure 12: Set DetectGPT, selected by LIME; $y_{i}=f(d_{i})=\text{human}$. Users only see explanations for the first document of each pair and annotate the second one
Figure 13: Set Solaiman, selected by Anchor; $y_{i}=f(d_{i})=\text{machine}$. Users only see explanations for the first document of each pair and annotate the second one
Appendix B Complementary results
DetectGPT0{0}250{250}f(x) = m0%single5%10%25%50%80%Tokens Masked0{0}250{250}f(x) = hGuoCounts0{0}250{250}f(x) = m0{0}250{250}f(x) = hSolaiman0{0}250{250}f(x) = m0{0}250{250}f(x) = hpadspaceremove
@@ Appendix B Complementary results
Detector Explainer Score p
Guo LIME 0.605 0.168
@@
Solaiman LIME 0.402 0.004
@@
DetectGPT Random 0.577 0.000
@@
Explainer Detector $\alpha$
SHAP Guo 0.896
Solaiman 0.812
DetectGPT 0.081
LIME Guo 0.478
Solaiman 0.439
DetectGPT 0.265
Anchor Guo 0.367
Solaiman 0.135
DetectGPT 0.129
Random Solaiman -0.137
Guo -0.137
DetectGPT -0.144
Table 6: Continuity: Per detector results
User Acc without User Acc with Change p
LIME Solaiman 0.681 0.569 -16.33% 0.057
LIME DetectGPT 0.792 0.681 -14.04% 0.134
LIME Guo 0.750 0.681 -9.26% 0.359
Anchor DetectGPT 0.806 0.750 -6.90% 0.289
SHAP DetectGPT 0.819 0.806 -1.69% 1.000
SHAP Solaiman 0.667 0.681 2.08% 1.000
Anchor Solaiman 0.569 0.583 2.44% 1.000
Anchor Guo 0.708 0.764 7.84% 0.344
SHAP Guo 0.778 0.847 8.93% 0.227
Table 7: Forward Simulation: Per group results.
DetectGPT Anchor Q1 2.15
Q2 2.04
Q3 2.11
LIME Q1 3.62
Q2 2.97
Q3 3.08
SHAP Q1 2.54
Q2 2.56
Q3 2.57
Solaiman Anchor Q1 3.17
Q2 3.18
Q3 3.15
LIME Q1 4.00
Q2 3.81
Q3 3.65
SHAP Q1 3.33
Q2 3.39
Q3 3.29
Guo Anchor Q1 2.39
Q2 2.22
Q3 2.26
LIME Q1 3.17
Q2 3.32
Q3 3.19
SHAP Q1 3.29
Q2 2.64
Q3 2.65
Table 8: Perceived Usefulness: Per group results. (3 = neutral, 5 = strongly agree)
@@
Table 9: Consistency: Per detector results
@@
Figure 16: Contrastivity: Fraction of tokens edited until $f(d_{i}^{\Omega})\neq f(d_{i})$
Appendix C Detector Benchmark Results
Acc F1 AUC TN FP FN TP ms/evaluation
DetectGPT GPT-2 @100 samples 0.502 0.000 0.500 153 0 152 0 15808
DetectGPT pythia-70m @100 samples 0.705 0.579 0.704 153 0 90 62 6391
DetectGPT pythia-70m (this paper) 0.744 0.664 0.743 150 3 75 77 898
Solaiman 0.921 0.922 0.921 139 14 10 142 19
Guo 0.990 0.990 0.990 153 0 3 149 18
Table 10: Performance of the detectors. Note that one could obtain better results for DetectGPT when using GPT-2 by adjusting the classification threshold.